# WildSAT: Learning Satellite Image Representations from Wildlife Observations

### v1

@@ Abstract
What does the presence of a species reveal about a geographic location? We posit that habitat, climate, and environmental preferences reflected in species distributions provide a rich source of supervision for learning satellite image representations. We introduce WildSAT, which pairs satellite images with millions of geo-tagged wildlife observations readily-available on citizen science platforms. WildSAT uses a contrastive learning framework to combine information from species distribution maps with text descriptions that capture habitat and range details, alongside satellite images, to train or fine-tune models. On a range of downstream satellite image recognition tasks, this significantly improves the performance of both randomly initialized models and pre-trained models from sources like ImageNet or specialized satellite image datasets. Additionally, the alignment with text enables zero-shot retrieval, allowing for search based on general descriptions of locations. We demonstrate that WildSAT achieves better representations than recent methods that utilize other forms of cross-modal supervision, such as aligning satellite images with ground images or wildlife photos. Finally, we analyze the impact of various design choices on downstream performance, highlighting the general applicability of our approach.
Figure 1: Wildlife observations can provide valuable supervision for learning satellite image representations. Known wildlife locations derived from human observations, coupled with descriptive information on species range, habitat, and other ecological attributes on Wikipedia, serve as a rich source of contextual information for satellite imagery. Our WildSAT approach leverages these additional data sources to (1) learn satellite image representations, and (2) enable zero-shot satellite image retrieval for identifying species habitats.
@@ 1 Introduction
Direct supervision in the form of paired images and labels (e.g. image tiles with labels denoting land cover type) can be prohibitively expensive to obtain at a global scale [27]. To address this, there is growing interest to develop methods that learn remote sensing representations from self-supervision [46, 44, 31], multiple paired modalities [60, 14, 45], or other auxiliary sources [68, 15]. For a supervision source to be useful, it needs to be globally distributed, correlated with the local landscape as viewed from an image, and able to discriminate regions at a fine spatial scale.
@@
One promising auxiliary supervision source is provided by the locations where different species of plants and animals can be found around the world. For example, at a coarser scale, Mountain Goats (Oreamnos americanus) are found in rugged mountainous areas, while at a finer scale, other species such as the Cactus Wren (Campylorhynchus brunneicapillus), are more habitat specialists as they are typically found in deserts nesting in spiny cacti (Figure 1). This information about where species can be found is an extremely rich source of potential supervision as it characterizes the local natural environment near each observation. It is also readily available online from citizen science platforms such as iNaturalist [2] and eBird [61] which contain hundreds of millions of wildlife location observations. While it has been shown that the locations of species can be used to improve fine-grained species classification [5, 9, 42, 60], it is not clear if this signal is useful for learning informative remote sensing representations.
@@
We introduce a new approach that uses signals derived from species location observations. We take inspiration from recent work that attempt to fuse multi-modal ecological data and remote sensing imagery into a shared common embedding space [59, 28, 60]. Using a contrastive learning objective, WildSAT minimizes the distance between embeddings that describe the same location, utilizing species observation data for combining the different modalities. Feature embeddings of satellite image, text, and location are aligned more closely when they originate from the same area, and moved further apart otherwise. Through this method, we utilize information about the preferred habitats of species to improve satellite image representations.
@@
We make the following contributions: (i) We introduce WildSAT, a new approach for learning remote sensing visual representations that exploits the supervision signal provided by the locations where different species have been observed. (ii) We demonstrate zero-shot satellite image retrieval capabilities using text descriptions of habitats and species. (iii) We present a detailed evaluation and show that WildSAT-derived representations outperform, and are complementary to, existing methods on a diverse set of models and downstream remote sensing tasks. (iv) We perform ablation studies to show the relative benefits of each component of our approach, and additionally show that WildSAT outperforms recent cross-modal learning approaches such as GRAFT [45] and TaxaBind [60].
@@ 2 Related Work
Supervised Learning. These methods leverage labels from tasks like object detection [73], instance segmentation [70], and classification [27, 62] in the satellite domain. Some works focus solely on image classification [73, 70, 72, 62, 52], while others learn from multiple label types. For example, SatlasPretrain [4] curated a large dataset with over 300 million labels across 137 categories, using domain experts, crowdsourced workers, and publicly available datasets (e.g. OpenStreetMap [23], WorldCover [66]). It is a unified model trained using multi-tasking across tasks (e.g. segmentation, object detection) and showed improved performance on various downstream remote-sensing tasks, significantly outperforming ImageNet-pretrained models and other baselines.
@@
Figure 2: Architecture for training and evaluating the satellite image encoder. (a) The training pipeline uses the location of a species, the satellite images at those locations, the environmental covariates, and the Wikipedia text associated with the species. In addition to the alignment of image, text, and location modalities, the encoder is encouraged to learn additional image features by using temporal and geometric image transformations on the input satellite image. (b) Downstream tasks use the frozen satellite image encoder with an additional trainable linear layer. Alternatively, the predicted image embeddings can be used for zero-shot retrieval via text queries.
Cross-Modal Learning. Recently, other works have explored adding other modalities while training on satellite images such as Synthetic Aperture Radar (SAR) [22], location [34, 59, 68], text [45, 28], and audio [60, 14, 33]. A common strategy is to use geo-tagged images and pre-trained image-text encoders such as CLIP [57]. Additional modalities are then projected to the embedding space of these existing models using a contrastive learning objective thereby aligning them [34, 45, 14, 68, 30]. Variations of this use the trained model for satellite image localization in GeoCLIP [68], bird species classification and mapping in BirdSAT [59], and improving image representation for plant species in CRISP [28]. Works such as GRAFT [45], TaxaBind [60], and GeoBind [14] align multiple modalities at the same time for cross-modal retrieval such as satellite to audio or ground image retrieval. Others explore methods outside of contrastive learning [64, 15] by employing supervised learning to fuse embeddings of different modalities such as satellite image with location in Sat-SINR [15] or text with location in LE-SINR [24] to predict species range maps, and fusing satellite images with bioclimatic rasters in SatBird [64] for bird species encounter rate prediction.
@@
Similar to those works, we investigate the use of species observation data to learn satellite image representations. Specifically, we explore how habitat preferences, climate, and other environmental factors encoded in species descriptions can serve as supervisory signals. While most previous work has focused on improving species distribution modeling [15, 64, 59, 41] or fine-grained image classification [60, 14, 43] using satellite images, our work improves satellite image representations using wildlife observations. Our experiments show that both randomly initialized models as well as strong baselines, such as Prithvi [31], SatasNet [4], and SeCo [46], benefit from this supervision on a wide range of satellite image recognition tasks (Table 1).
@@ 3 Method
We define the problem as follows: given an image encoder $f_{\theta}:\mathbf{I}\rightarrow\mathbf{z}$ with parameters $\theta$, we want to find an optimal set of parameters $\theta^{*}$ that improves the performance of $f$ on various remote sensing tasks through a robust satellite image feature representation $\mathbf{z}$. It takes an image $\mathbf{I}\in\mathbb{R}^{W\times H\times 3}$ as input and outputs an embedding $\mathbf{z}\in\mathbb{R}^{d}$. We propose to optimize $\theta$ using our WildSAT framework using data consisting of satellite images, locations, environmental covariates, and text. We hypothesize that leveraging known environmental context around each species observation allows for more effective optimization of model parameters (Figure 1).
@@ 3.1 WildSAT
To supplement satellite images, we take advantage of additional modalities that naturally align based on the distribution of species throughout the globe. Information on the habitat of species can provide a rich source of supervision for improving satellite image representations, and we describe how to leverage this through our proposed framework WildSAT. Figure 2 shows the overall architecture used to train a satellite image encoder $f_{\theta}$. The encoder $f$ can be any architecture (e.g. a ResNet50 [25], ViT-B/16 [16], etc.). The initial parameters $\theta$ can be randomly initialized, pre-trained on a different domain (e.g. ImageNet [13]), or pre-trained on a related dataset (e.g. SatlasPretrain [4]). The output embedding $\mathbf{z}$ can be used for downstream remote sensing tasks such as classification and zero-shot image retrieval.
@@
Location encoder. SINR [11] is primarily used for predicting the presence and absence of species at a location. It does this by training on large collections of species observation data. It takes location $\mathbf{a}_{i}=(lat,lon)\in\mathbb{R}^{2}$ and, optionally, the corresponding environmental covariates $\text{env}(\mathbf{a}_{i})\in\mathbb{R}^{20}$ to produce a location embedding in $\mathbb{R}^{256}$.
@@
Large Language Model. GritLM [50] is a large language model (LLM) that outputs a fixed-length embedding in $\mathbb{R}^{4096}$ given a text input. It is trained to handle text of arbitrary length, making it suitable for the varying lengths of text obtained from sources such as Wikipedia [3].
@@
Satellite Image Encoder. Given an initial image encoder $f$, we add three sets of linear layers to predict embeddings for images ($\mathbf{z}_{I^{t}}$), text ($\mathbf{z}_{\text{txt}}$), and locations ($\mathbf{z}_{\text{loc}}$). Similarly, both the pre-trained LLM (GritLM) and location encoder (SINR) have an added trainable linear layer to project their respective feature embeddings to $\mathbb{R}^{d}$ as $\mathbf{e}_{\text{txt}}$ and $\mathbf{e}_{\text{loc}}$, respectively. In addition to text and location, we also fine-tune the model on image-specific features by forcing embeddings of similar satellite images to be close to one another. For two satellite images $\mathbf{I}^{t_{1}}$ and $\mathbf{I}^{t_{2}}$ taken at the same location but at different times, their corresponding feature representations should be similar. We also apply geometric augmentations $T$ such as flipping and random cropping on the latter image such that $f(\mathbf{I}^{t_{1}})\approx f(T(\mathbf{I}^{t_{2}}))$. Doing this encourages the model to learn meaningful image features by distinguishing between similar/dissimilar images while refining the same features through the text, location, and environment.
3.2 Dataset
To train the model, we combine images, text, location, and environmental covariates from publicly available datasets [11, 24, 19, 3, 17]. For a given species, we obtain its corresponding observation data through iNaturalist [67], and a text description of its preferred habitat from its corresponding Wikipedia [3] page (Figure 1). Satellite images are then retrieved based on the species observation locations. We describe each component of the dataset below.
Location Observation Data. iNaturalist [67] observations consist of latitude and longitude values denoting the locations where a species has been observed. We use the dataset from SINR [11] that contains 35.5 million observations of 47,375 different species. It is composed of $\{(\mathbf{a}_{i},b_{i})\}_{i=1}^{N}$ pairs where $\mathbf{a}_{i}=(lat,lon)$ is the location and $b_{i}\in\{s_{1},s_{2},\dots,s_{M}\}$ is an integer encoding the species name of the observed species. The locations of each species are used as the basis for matching other sources of data such as environmental covariates, satellite images, and text.
Environmental Covariates. Environmental covariates are obtained from WorldClim2 [19] for a location $\mathbf{a}_{i}$. The data aggregates temperature and precipitation values and are often used for ecological applications. It returns a vector for each location $\text{env}(\mathbf{a}_{i})\in\mathbb{R}^{20}$. We use five minute resolution data and bilinearly interpolate between data points to get values for specific locations.
Satellite Images. Sentinel-2 satellite images are used with a resolution of ten meters per pixel, and size 512$\times$512. The satellite provides 13 spectral bands at different resolutions and has a revisit frequency of five days near the equator. For simplicity and general applicability, since different satellites could sample different sets of bands, we only use RGB bands. We additionally remove satellite images with significant cloud cover. Satellite images taken at the same locations but at a different time are collected as a form of augmentation for training. A total of 305,689 satellite images are collected from EU’s Sentinel data [17, 4].
Text. For each species, we use readily available text data. The text is taken from Wikipedia [3] as in LE-SINR [24]. The corresponding page typically has several sections describing different aspects of the species such as its habitat, range, behavior, taxonomy, etc. Each section is processed separately such that one text embedding corresponds to a single section of the Wikipedia text. Similar to [24], we remove sections that do not provide information about the species (e.g. references, bibliography, links). The final text dataset contains 127,484 sections from 37,889 species.
For a single satellite image-species location match, there could be multiple text embeddings that correspond to the multiple sections of the species Wikipedia [3] page. Taking these into account, there are a total of 980,376 training samples with location, satellite image, and text. During training, we randomly sample one section of text for every satellite image-species location match, resulting in effectively 134,890 iterations per epoch. We show the spatial distribution of the data in the Appendix.
@@ 3.3 Training
Our framework uses a contrastive learning objective to improve the satellite image encoder embeddings. We jointly optimize the parameters of the model $f_{\theta}$ and the additional linear layers through the training objective in Eqn. 1. These loss terms correspond to a contrastive objective over image embeddings ($\mathcal{L}_{\text{img}}$), text embeddings ($\mathcal{L}_{\text{txt}}$), and location embeddings ($\mathcal{L}_{\text{loc}}$) of $f$, and the overall objective is:
@@
We compute the distance between two sets of embeddings $\mathbf{Z}$ and $\mathbf{E}$ using a minibatch of $n$ samples with the $i$-th embedding in $\mathbf{Z}$ aligned with the $i$-th embedding of $E$.s That is, given two sets of embeddings $\mathbf{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n}\}$ and $\mathbf{E}=\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$, the embedding $\mathbf{z}_{i}$ is matched to $\mathbf{e}_{i}$ against other embeddings $\mathbf{e}_{1,\ldots,n}$ using the loss in Eqn. 2.
@@
Based on the InfoNCE loss [53, 57], Eqn. 3 matches embedding $\mathbf{z}_{i}$ with the corresponding embedding $\mathbf{e}_{i}$ by minimizing the distance between them with respect to the other embeddings in $\mathbf{E}$, with temperature hyperparameter $\tau$.

### v2

@@ Abstract
Species distributions encode valuable ecological and environmental information, yet their potential for guiding representation learning in remote sensing remains underexplored. We introduce WildSAT, which pairs satellite images with millions of geo-tagged wildlife observations readily-available on citizen science platforms. WildSAT employs a contrastive learning approach that jointly leverages satellite images, species occurrence maps, and textual habitat descriptions to train or fine-tune models. This approach significantly improves performance on diverse satellite image recognition tasks, outperforming both ImageNet-pretrained models and satellite-specific baselines. Additionally, by aligning visual and textual information, WildSAT enables zero-shot retrieval, allowing users to search geographic locations based on textual descriptions. WildSAT surpasses recent cross-modal learning methods, including approaches that align satellite images with ground imagery or wildlife photos, demonstrating the advantages of our approach. Finally, we analyze the impact of key design choices and highlight the broad applicability of WildSAT to remote sensing and biodiversity monitoring.
Figure 1: Wildlife observations can provide valuable supervision for learning satellite image representations. Known wildlife locations derived from human observations, coupled with descriptive information on species range, habitat, and other ecological attributes on Wikipedia, serve as a rich source of contextual information for satellite imagery. Our WildSAT approach leverages these additional data sources to (i) learn robust satellite image representations for downstream tasks, and (ii) complement and further improve existing models using continual pre-training.
@@ 1 Introduction
Direct supervision in the form of paired images and labels (e.g. image tiles with labels denoting land cover type) can be prohibitively expensive to obtain at a global scale [28]. To address this, there is growing interest to develop methods that learn remote sensing representations from self-supervision [48, 46, 32], multiple paired modalities [63, 15, 47], or other auxiliary sources [72, 16]. A useful supervision source needs to be globally distributed, correlated with the local landscape as viewed from an image, and able to discriminate regions at a fine spatial scale.
@@
A promising auxiliary supervision source is provided by locations where different species of plants and animals can be found around the world. For example, Mountain Goats (Oreamnos americanus) are found in rugged mountainous areas, while habitat specialists like the Cactus Wren (Campylorhynchus brunneicapillus) are typically found in deserts nesting in spiny cacti (Fig. 1). Species location data offer a rich source of supervision, reflecting the local natural environment around each observation. It is also readily available from citizen science platforms such as iNaturalist [2] and eBird [64] which host hundreds of millions of wildlife observations. While species location data has improved fine-grained species classification [6, 10, 44], its potential for learning remote sensing representations remains unclear. Prior works have largely relied on anthropogenic labels (e.g. human-made features like roads, buildings, industrial areas) to learn satellite image representations [69, 47, 41], whereas we explore using wildlife observations as a complementary and potentially valuable signal.
@@
We introduce a new approach that uses signals derived from species location observations. We take inspiration from recent work that attempt to fuse multi-modal ecological data and remote sensing imagery into a shared common embedding space [62, 29, 63]. WildSAT uses a contrastive learning objective to align satellite image, text, and location based on species observation data, bringing embeddings from the same area closer together and pushing those from different areas further apart. Through this method, we utilize information about the preferred habitats of species to improve satellite image representations.
@@
We make the following contributions: (i) We introduce WildSAT, a new approach to learning remote sensing representations using species observation locations as a supervisory signal. (ii) We show WildSAT-derived representations are competitive with state-of-the-art satellite representations, while enabling zero-shot satellite image retrieval. (iii) We present a thorough evaluation that highlights WildSAT-derived representations not only outperform but also complement existing methods focused on anthropogenic labels by incorporating wildlife information. (iv) We perform ablation studies to show the impact of each component of our approach, and show WildSAT outperforms recent cross-modal methods like GRAFT [47] and TaxaBind [63]. The code and dataset are available at https://github.com/cvl-umass/wildsat.
@@ 2 Related Work
Supervised learning. These methods leverage labels from tasks like object detection [77], instance segmentation [74], and classification [28, 65] in the satellite domain. Some works focus solely on image classification [77, 74, 76, 65, 54], while others learn from multiple label types. For example, SatlasPretrain [5] curated a large dataset with over 300 million labels across 137 categories, using domain experts, crowdsourced workers, and publicly available datasets (e.g. OpenStreetMap [24], WorldCover [70]). It is a unified model jointly trained for multiple tasks (e.g. segmentation, object detection) and showed improved performance on various downstream remote-sensing tasks, significantly outperforming ImageNet-pretrained models and other baselines.
@@
Figure 2: Architecture for training and evaluating the satellite image encoder. (a) The training pipeline uses the location of a species, the satellite images at those locations, the environmental covariates, and the Wikipedia text associated with the species. In addition to the alignment of image, text, and location modalities, the encoder is encouraged to learn additional image features by using temporal and geometric image transformations on the input satellite image. (b) Downstream tasks use the frozen satellite image encoder with an additional trainable layer (or layers). Alternatively, the predicted image embeddings can be used for zero-shot retrieval via text queries.
Cross-modal learning. Recently, other works have explored adding other modalities while training on satellite images [23, 35, 62, 72, 47, 29, 63, 15, 34]. A common approach uses geo-tagged images and pre-trained image-text encoders like CLIP [59], aligning new modalities to their embedding space using contrastive learning [35, 47, 15, 72, 31, 41]. This strategy has been used for various tasks: satellite image localization in GeoCLIP [72], bird species classification and mapping in BirdSAT [62], and improving plant species image representations in CRISP [29]. Models like GRAFT [47], TaxaBind [63], RemoteCLIP [41], and GeoBind [15] align multiple modalities at the same time for cross-modal retrieval and zero-shot tasks. Zermatten et al. [81] have also demonstrated the benefits of aligning satellite imagery with species observation data for zero-shot classification. TaxaBind was the first to use species geographic locations and satellite imagery, but it focuses on ecological tasks rather than satellite image tasks. We also expand on their approach by leveraging open-source Wikipedia text instead of taxonomic hierarchy data, offering a more diverse supervision for satellite image representations. Beyond contrastive learning, other methods use supervised learning to fuse embeddings of different modalities for predicting species range maps and encounter rates [16, 25, 67]. Most similar to our work is WikiSatNet [69] which uses location-aligned Wikipedia articles and satellite images to improve satellite image representations. However, while WikiSatNet and previous works [47, 69, 41, 35] primarily focus on anthropogenic data, we explore the impact of wildlife observations. Specifically, we investigate how species distributions—capturing habitat preferences, climate, and environmental factors—can serve as powerful signals.
@@
While previous work has focused on improving species distribution modeling [16, 67, 62, 43] or fine-grained image classification [63, 15, 45] using satellite images, our work improves satellite image representations using wildlife observations. Our experiments show that both randomly initialized models and strong baselines, such as Prithvi [32], SatlasNet [5], and SeCo [48], benefit from this supervision on a wide range of satellite image tasks (Fig. 3, Tab. 1).
@@ 3 Method
We define the problem as follows: given an image encoder $f_{\theta}:\mathbf{I}\rightarrow\mathbf{z}$ with parameters $\theta$, we want to find an optimal set of parameters $\theta^{*}$ that improves the performance of $f$ on various remote sensing tasks through a robust satellite image feature representation $\mathbf{z}$. It takes an image $\mathbf{I}\in\mathbb{R}^{W\times H\times 3}$ as input and outputs an embedding $\mathbf{z}\in\mathbb{R}^{d}$. We propose to optimize $\theta$ using our WildSAT framework using data consisting of satellite images, locations, environmental covariates, and text. We hypothesize that leveraging known environmental context around each species observation (Fig. 1) allows for more effective optimization of model parameters.
@@ 3.1 WildSAT
To supplement satellite images, we take advantage of additional modalities that naturally align based on the distribution of species throughout the globe. Information on species habitat can provide a rich source of supervision for improving satellite image representations, and we describe how to leverage this through our proposed WildSAT framework. Fig. 2 shows the overall architecture used to train a satellite image encoder $f_{\theta}$. The encoder $f$ can be any architecture (e.g. a ResNet50 [26], ViT-B/16 [17], etc.). The initial parameters $\theta$ can be randomly initialized, pre-trained on a different domain (e.g. ImageNet [14]), or pre-trained on a related dataset (e.g. SatlasPretrain [5]). The output embedding $\mathbf{z}$ can be used for downstream remote sensing tasks such as classification and zero-shot image retrieval.
@@
Location encoder. SINR [12] is mainly used to predict the presence and absence of species at a location by training on large collections of species observation data. It takes location $\mathbf{a}_{i}=(lat,lon)\in\mathbb{R}^{2}$ and, optionally, the corresponding environmental covariates (e.g. weather and climate data) to produce a location embedding in $\mathbb{R}^{256}$.
@@
Large language model. GritLM [52] is a large language model (LLM) that outputs a fixed-length embedding in $\mathbb{R}^{4096}$ given a text input. It is trained to handle text of arbitrary length, making it suitable for longer and varying lengths of text obtained from sources such as Wikipedia [4].
@@
Satellite image encoder. Given an initial image encoder $f$, we add three sets of linear layers to predict embeddings for images ($\mathbf{z}_{I^{t}}$), text ($\mathbf{z}_{\text{txt}}$), and locations ($\mathbf{z}_{\text{loc}}$). Similarly, both the pre-trained LLM (GritLM) and location encoder (SINR) have an added trainable linear layer to project their respective feature embeddings to $\mathbb{R}^{d}$ as $\mathbf{e}_{\text{txt}}$ and $\mathbf{e}_{\text{loc}}$, respectively. In addition to text and location, we also fine-tune the model on image-specific features by forcing embeddings of similar satellite images to be close to each another. For two satellite images $\mathbf{I}^{t_{1}}$ and $\mathbf{I}^{t_{2}}$ taken at the same location but at different times, their corresponding feature representations should be similar. We also apply geometric augmentations $T$ such as flipping and random cropping on the latter image such that $f(\mathbf{I}^{t_{1}})\approx f(T(\mathbf{I}^{t_{2}}))$. Doing this encourages the model to learn meaningful image features by distinguishing between similar and dissimilar images while refining the same features through text, location, and environment.
@@ 3.2 Training
Our framework uses a contrastive learning objective to improve satellite image encoder embeddings. We jointly optimize the parameters of the model $f_{\theta}$ and the additional linear layers through the training objective in Eqn. 1. These loss terms correspond to a contrastive objective over image embeddings ($\mathcal{L}_{\text{img}}$), text embeddings ($\mathcal{L}_{\text{txt}}$), and location embeddings ($\mathcal{L}_{\text{loc}}$) of $f$. The embeddings $\mathbf{z}_{\mathbf{I}^{t}},\mathbf{z}_{\text{txt}},\mathbf{z}_{\text{loc}}$ are linear projections of the image embedding for each of the modalities (see Fig. 2). Eqn 1 shows the objective.
@@
We compute the distance between two sets of embeddings $\mathbf{Z}$ and $\mathbf{E}$ using a minibatch of $n$ samples with the $i$-th embedding in $\mathbf{Z}$ aligned with the $i$-th embedding of $\mathbf{E}$. That is, given two sets of embeddings $\mathbf{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n}\}$ and $\mathbf{E}=\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$, the embedding $\mathbf{z}_{i}$ is matched to $\mathbf{e}_{i}$ against other embeddings $\mathbf{e}_{1,\ldots,n}$ using the loss in Eqn. 2.
@@
Based on InfoNCE loss [55, 59], Eqn. 3 matches the embedding $\mathbf{z}_{i}$ with the corresponding embedding $\mathbf{e}_{i}$ by minimizing the distance between them with respect to the other embeddings in $\mathbf{E}$, with temperature hyperparameter $\tau$.